# Topologically protected Bell-cat states in a simple spin model

### v2

@@ III.3 Replacing the identical spins with a single bosonic mode
Coherent cat states of light have been achieved using a variety of methods, so it is important to determine whether it is possible to do the same using our protocol. To this end, we replace the identical spin operators in Eq. (1) with operators of a single bosonic mode, $2\hat{S}_{-}/N\to\hat{a}$ and $2\hat{S}_{+}/N\to\hat{a}^{\dagger}$, where $\hat{a}^{\dagger}$ ($\hat{a}$) is the creation (annihilation) operator of the mode obeying the usual commutation relation $[\hat{a},\hat{a}^{\dagger}]=1$. The Hamiltonian becomes
$\hat{H}\to\hat{H}_{a}=v\hat{\sigma}_{x}+w\left(\hat{a}^{\dagger}\hat{\sigma}_{-}+\hat{a}\hat{\sigma}_{+}\right).$ (12)
@@
$\left[\left(v+\frac{w}{\sqrt{2}}x\right)\hat{\sigma}_{x}+i\frac{w}{\sqrt{2}}\frac{\partial}{\partial x}\hat{\sigma}_{y}\right]\tilde{\chi}_{\uparrow(\downarrow)}(x)=0\$ (14)
@@
Equation 17 shows that when both $w$ and $v$ are nonzero, the winding number changes once at $|\alpha|=v/w$, which corresponds to the location of the bound state. Figure 5 displays the Fock space probability distributions of the bound state, obtained numerically by diagonalizing Eq. (16) (red dashed curve), alongside the coherent state centered at $m_{a}=|\alpha|^{2}=|v/w|^{2}$ (black solid curve). The excellent agreement between the two curves confirms that the bound state is well-approximated by a coherent state localized at this point. The inset shows the energy spectrum of $\hat{H}_{a}$, highlighting two bound states at $E=0$ in red. The first is the bound state featured in the main figure, while the second arises due to the truncation of the bosonic mode’s Hilbert space at $N_{a}=201$, which introduces an artificial boundary. According to the bulk-boundary correspondence, this boundary supports an additional bound state at the opposite edge. This degeneracy can be lifted—even in the truncated Hilbert space—by adding a perturbative term to $\hat{H}_{a}$ proportional to the mode energy, $\hat{a}^{\dagger}\hat{a}$. The perturbation must remain small enough to avoid closing the energy gap between the bound state and the bulk, which would otherwise destroy the true bound state.
Figure 5: Probability distribution of the bound state (red, dashed) of $\hat{H}_{a}$ in Eq. (16) alongisde a coherent state (black, solid). The number of states of the bosonic mode is truncated at $N_{a}=201$ and the parameters are $v=7$ and $w=1$, so both states are centered at $m_{a}=49$. Inset: Zoom in of the spectrum of $\hat{H}_{a}$ around the bound states (red) at $E=0$.
@@ IV Conclusions and Discussion
We have demonstrated that in a CS model obeying chiral symmetry, a topological phase transition takes place. In the topologically nontrivial phase, a pair of protected bound states emerges within the Fock space of the identical spins surrounding the CS. Each member of this pair is coupled to a different state of the CS. Notably, these bound states exhibit a unique characteristic — their positions in Fock space can be adjusted by tuning parameters in the system, distinguishing them from typical bound states found in many condensed matter systems. This distinction stems from the spherical geometry of the Hilbert space of identical spins, which differs from the circular and toroidal geometry found in the quasi-momentum of 1D and 2D periodic lattices, respectively.
@@
Although the version of the CS model we have analyzed is quite simple, it can be expanded to be more general because the main criterion needed to generate a BC state is the presence of mobile topologically protected bound states. This requirement means that the Hamiltonian must obey chiral symmetry which leaves room for additional terms of the form $\hat{\sigma}_{x}f(\hat{S}_{x},\hat{S}_{y},\hat{S}_{z})$ and $\hat{\sigma}_{y}g(\hat{S}_{x},\hat{S}_{y},\hat{S}_{z})$ where $f$ and $g$ denote general functions of the identical spins’ operators. The zero energy eigenstates (there may be many pairs of such states depending on the functions) will also be eigenstates of the chiral symmetry operator, $\hat{\sigma}_{z}$, and will serve as the target BC state(s) in similar driving protocols to the one considered in this paper. In experiments, it may be impossible to neglect chiral symmetry-breaking terms. Two such terms that may appear in the CS model are $v_{z}\hat{\sigma}_{z}$ and $w_{z}\hat{\sigma}_{z}\hat{S}_{z}$. Separately they destroy the bound states, however, together it may be possible to tune the parameters of the system such that their contributions cancel for a given bound state, $v_{z}\hat{\sigma}_{z}+w_{z}\hat{\sigma}_{z}\langle\hat{S}_{z}\rangle_{B}=0$. Although further analysis is necessary to explore this possibility, it could potentially offer a workaround for certain persistent symmetry-breaking terms.
@@ Acknowledgements.
Appendix A Target eigenstate wave functions in the topologically trivial phase
@@
where we have used the relation $n=\frac{N}{2}\cos\theta$. The overall sign signifies whether the CS is in an even ($+$) or odd (--) superposition and corresponds to the top and bottom energy bands, respectively, in Figure 4 (a). In the trivial phase, the even parity bound state is the ground state of the top band which we know is centered at $(n,\phi)=(0,\pi)$, so we expand the energy around those points keeping up to quadratic order in both variables
@@
In the semiclassical theory, $n$ and $\phi$ are conjugate variables obeying the commutation relation $[n,\phi]=-i$, so we can make the transformation $\phi\to i\frac{d}{dn}$ resulting in

### v3

@@ III.3 Replacing the identical spins with a single bosonic mode
Coherent cat states of light have been achieved using a variety of methods, so it is important to determine whether it is possible to do the same using our protocol. To this end, we replace the identical spin operators in Eq. (1) with operators of a single bosonic mode, $2\hat{S}_{-}/N\to\hat{a}$ and $2\hat{S}_{+}/N\to\hat{a}^{\dagger}$, where $\hat{a}^{\dagger}$ ($\hat{a}$) is the creation (annihilation) operator of the mode obeying the usual commutation relation $\hat{a},\hat{a}^{\dagger}]=1$. The Hamiltonian becomes
[…]
@@
$\left[\left(v+\frac{w}{\sqrt{2}}x\right)\hat{\sigma}_{x}+i\frac{w}{\sqrt{2}}\frac{\partial}{\partial x}\hat{\sigma}_{y}\right]\tilde{\chi}_{\uparrow(\downarrow)}(x)=0\\$ (18)
@@
Equation 21 shows that when both $w$ and $v$ are nonzero, the winding number changes once at $|\alpha|=v/w$, which corresponds to the location of the bound state. Figure 5 displays the Fock space probability distributions of the bound state, obtained numerically by diagonalizing Eq. (20) (red dashed curve), alongside the coherent state centered at $m_{a}=|\alpha|^{2}=|v/w|^{2}$ (black solid curve). The inset shows the energy spectrum of $\hat{H}_{a}$, highlighting two bound states at $E=0$ in red. The first is the bound state featured in the main figure, while the second arises due to the truncation of the bosonic mode’s Hilbert space at $N_{a}=201$, which introduces an artificial boundary. According to the bulk-boundary correspondence, this boundary supports an additional bound state at the opposite edge. This degeneracy can be lifted, even in the truncated Hilbert space, by adding a perturbative term to $\hat{H}_{a}$ proportional to the mode energy, $\hat{a}^{\dagger}\hat{a}$. The perturbation must remain small enough to avoid closing the energy gap between the bound state and the bulk, which would otherwise destroy the true bound state.
Figure 5: Probability distribution of the bound state (red, dashed) of $\hat{H}_{a}$ in Eq. (20) alongside a coherent state (black, solid). The number of states of the bosonic mode is truncated at $N_{a}=201$ and the parameters are $v=7$ and $w=1$, so both states are centered at $m_{a}=49$. Inset: Zoom in of the spectrum of $\hat{H}_{a}$ around the bound states (red) at $E=0$.
@@ IV Conclusions and Discussion
We have demonstrated that in a CS model obeying chiral symmetry, a topological phase transition takes place. In the topologically nontrivial phase, a pair of protected bound states emerges within the Fock space of the identical spins surrounding the CS. Each member of this pair is coupled to a different state of the CS. Notably, these bound states exhibit a unique characteristic: their positions in Fock space can be adjusted by tuning parameters in the system, distinguishing them from typical bound states found in many condensed matter systems. This distinction stems from the spherical geometry of the Hilbert space of identical spins, which differs from the circular and toroidal geometry found in the quasi-momentum of 1D and 2D periodic lattices, respectively.
@@
Although the version of the CS model we have analyzed is quite simple, it can be expanded to be more general because the main criterion needed to generate a BC state is the presence of mobile topologically protected bound states. This requirement means that the Hamiltonian must obey chiral symmetry which leaves room for additional terms of the form $\hat{\sigma}_{x}f(\hat{S}_{x},\hat{S}_{y},\hat{S}_{z})$ and $\hat{\sigma}_{y}g(\hat{S}_{x},\hat{S}_{y},\hat{S}_{z})$ where $f$ and $g$ denote general functions of the identical spins’ operators. The zero energy eigenstates (there may be many pairs of such states depending on the functions) will also be eigenstates of the chiral symmetry operator, $\hat{\sigma}_{z}$, and will serve as the target BC state(s) in similar driving protocols to the one considered in this paper. In experiments, it may be impossible to neglect chiral symmetry breaking terms. We mentioned two such terms that may appear in the CS model which were $v_{z}\hat{\sigma}_{z}$ and $2w_{z}\hat{\sigma}_{z}\hat{S}_{z}/N$. We considered them as perturbations, but it may be possible to mitigate their effects at arbitrary strengths. Separately they destroy the bound states, however, together it may be possible to tune the parameters of the system such that their contributions cancel for a given bound state, $v_{z}\hat{\sigma}_{z}+w_{z}\hat{\sigma}_{z}\langle\hat{S}_{z}\rangle_{B}=0$. Although further analysis is necessary to explore this possibility, it could potentially offer a workaround for certain persistent symmetry breaking terms. Finally, throughout the paper we assume that all surrounding spins couple equally to the central spin. We remark that both chiral and parity symmetries survive if this condition is relaxed, so each spin sector still hosts BC states. If the variation in coupling strengths is small enough that the system’s dynamics remain mostly confined to the spin-$N/2$ sector, the BC state can still be generated.
@@ Acknowledgements.
Appendix A Finite Size Scaling
Here, we present calculations of the finite size scaling of both $\gamma^{\prime}$ and the fidelity between the initial coherent state and the target eigenenergy state which becomes the BC state in the topological phase. We relax the minimization condition over the values of $v$ in Eq. (15), so that we can calculate the finite size scaling of $\gamma^{\prime}$ in the trivial phase, the topological phase and at the critical point. Figure 7 shows a plot of $\ln(\gamma^{\prime})$ as a function of $\ln(1/N)$ for $v=0.1w$ (topological phase, red circles), $v=w$ (critical point, black triangles) and $v=2w$ (trivial phase, blue squares). The best fit lines displayed in the figure indicate scaling exponents consistent with $N^{-1}$ in both the topological and trivial phases, and $N^{-4/3}$ at the critical point. From the image we also see that $\gamma^{\prime}$ is the smallest at the critical point which is expected as the energy gap is smallest there.
Figure 8 shows the fidelity between the initial coherent state and the target state as a function of $1/N$. The best fit line takes the form $y=c_{0}+c_{1}x$, indicating that the system size dependent term scales as $N^{-1}$. Although we have calculated both constants numerically, as shown in the image, $c_{0}$ can be obtained analytically by evaluating the fidelity between the expressions for the initial coherent state and the target state in Eqns. (11) and (30), respectively. This gives
$c_{0}=\frac{2\left(1-\frac{w}{v}\right)^{1/4}}{1+\sqrt{1-\frac{w}{v}}}$ (26)
and even for moderate initial values of $v$ such as $v_{0}=2w$, we get $c_{0}=0.98517$, which agrees well with the numerical result shown in the image.
Figure 7: Plot of $\ln(\gamma^{\prime})$ as a function of $\ln(1/N)$. The data corresponds to $v=0.1w$ (topological phase, red circles), $v=w$ (critical point, black triangles) and $v=2w$ (trivial phase, blue squares) while the equations are best fit lines. The range of system sizes is $1000\leq N\leq 2000$ and $w=1$.
Figure 8: Fidelity of the initial coherent state in Eq. (10) and the target state as a function of $1/N$. The range of system sizes is $100\leq N\leq 500$ and the parameters are $v_{0}=2w$ and $w=1$.
Appendix B Target eigenstate wave functions in the topologically trivial phase
@@
where we have used the relation $n=\frac{N}{2}\cos\theta$. The overall sign signifies whether the CS is in an even (++) or odd (-) superposition and corresponds to the top and bottom energy bands, respectively, in Figure 4 (a). In the trivial phase, the even parity bound state is the ground state of the top band which we know is centered at $(n,\phi)=(0,\pi)$, so we expand the energy around those points keeping up to quadratic order in both variables
@@
In the semiclassical theory, $n$ and $\phi$ are conjugate variables obeying the commutation relation $n,\phi]=-i$, so we can make the transformation $\phi\to i\frac{d}{dn}$ resulting in